# Topological Kondo semimetals emulated in hetero-bilayer transition metal dichalcogenides

### v1

@@ I Introduction
Figure 1: (a) Global phase diagram of Kondo lattice models at commensurate filling factor with ($\nu_{c}\approx\nu_{d}\approx 1$), where $c$ and $d$ correspond to light and heavy orbitals, respectively. Vertical axis $G$ stands for the frustration strength, and horizontal axis $\rho J_{K}$ describes the Kondo coupling. The blue and black lines correspond to Kondo-destruction (dehybridization) and magnetic transitions, respectively. They separate the parameter region into four phases: (i) $M_{S}$, a magnetic ordered phase with a small Fermi surface; (ii) $P_{S}$, a paramagnetic phase with a small Fermi surface; (iii) $M_{L}$, a magnetic ordered phase with a large Fermi surface; and (iv) Kondo-screened phase, which is either a Kondo semimetal (KSM) or a Kondo insulator (KI). The dashed lines describe distinct sequences of quantum phase transitions between the Kondo-destroyed phases and the Kondo-screened ones. Adapted from Ref. [33]. (b) The lattice structure of the $2H$ phase of single layer transition metal dichalcogenides. The purple and yellow orbs represent the metal and chalcogen atoms, respectively. This lattice structure already breaks the inversion symmetry. Lattice structure data retrieved from the Materials Project for Te2Mo (mp-1023935) from database version v2023.11.1 [34] (c) Spin-valley locking in the TMDC materials. $\mathcal{K},\mathcal{K}^{\prime}$ denote the two valleys in the TMDC system. Due to the AB-stacking structure, the spin polarizations from the two TMDC layers in the same valley are different, thereby reducing the interlayer hybridization. (d) Noninteracting band structure of the two-orbital effective tight binding model with $t_{c}=9\leavevmode\nobreak\ \rm meV$, $t_{d}=4.5\leavevmode\nobreak\ \rm meV$, $t_{cd}=1.5\leavevmode\nobreak\ \rm meV$ and $\epsilon_{D}=0\leavevmode\nobreak\ \rm meV$, which well-captures the properties of the active moiré bands in AB-stacked $\rm MoTe_{2}/WSe_{2}$.
@@ I.2 Kondo effect and its realization in heterobilayer TMDCs
Kondo lattice systems [42] represent a canonical setting to realize a landscape of correlated quantum phases. The contrast in the width of the $f$-electron band and that of the $spd$-electron bands causes orbital-selective Mott correlations [43, 44]. In particular, metallic quantum criticality and strange metallicity have been the main stay of the modern era of this field, fueled by the notion of Kondo destruction [45, 46, 47]. This notion captures the dynamical competition between the Kondo and Ruderman–Kittel–Kasuya–Yosida (RKKY) interactions, leading, for example, to an interacting quantum critical point with dynamical “Planckian” scaling [45]. The Kondo destruction notion is supported by a large body of experiments in a variety of quantum critical heavy fermion metals  [48, 49, 50, 51, 52, 53, 54]. When combined with geometrical frustration and related quantum fluctuations of the local moments, this notion leads to a global phase diagram [55, 56, 57, 58], delineating the interplay between the local-moment magnetism and paramagnetism, the “large” Fermi surface associated with the Kondo-screened ground state, and the “small” Fermi surface associated with the destruction of the Kondo effect [59, 60].
@@
In other words, if we compare the energy scale in the AB-stacked heterobilayer TMDC system, it will fall into the parameter regime with $t_{c}>t_{d}>\tilde{t}_{cd}$ (where $\tilde{t}_{cd}\approx\pi\rho_{0,c}t^{2}_{cd}$ with $\rho_{0,c}$ being the typical bare density of states of the $c$ band). This is different from what happens in the canonical heavy fermion systems [42, 43, 44] or in the effective Kondo lattice models realized in geometrically-enabled-flat-band systems [61, 62, 63], where $t_{c}>\tilde{t}_{cd}>t_{d}$.
This “inverted” hierarchy of energy scales has important consequences. One concerns the coherent temperature scale, which characterizes the crossover between the high-temperature incoherent scattering regime and the low temperature Fermi liquid regime. Due to the inverted energy scale hierarchy, the coherence temperature relevant to the strongly correlated metal region depends not only on the hybridization but also on the contribution from the intra-orbital kinetic energy of the heavy electron itself [9]. Accordingly, this coherent temperature scale will be unexpectedly large, as indeed observed experimentally [3]. As we will show, this energy scale hierarchy plays an important role in realizing a topological Kondo semimetal phase in the commensurate filling that we will study.
@@ I.3 Topological Kondo lattice Systems at commensurate fillings: Topological Kondo semimetals
In recent years, topological Kondo lattice models [35, 36, 37] and materials [38, 39] have been advanced, which realize the cooperation between the inherent topology of the noninteracting bands and the correlation physics associated with the Kondo effect at commensurate electron fillings. At such fillings, a Kondo insulator gap would have been opened; its prevention by the the lattice symmetry that underlies the band topology gives rise to Weyl Kondo semimetals.
@@ I.4 Scope of the present work
Recently, the Kondo effect and its destruction have been theoretically discussed in models for the AB-stacked $\rm{MoTe_{2}/WSe_{2}}$ [8, 9]. The focus was at incommensurate fillings. The inherently topological nature of the electronic bands raises the prospect that this moiré structure can emulate topological Kondo lattice systems. It is the objective of this work to explore this prospect.
Our focus is on the case with a fixed moiré unit cell filling $\nu_{\rm tot}=\nu_{\rm Mo}+\nu_{\rm W}=1+1$ in the AB-stacked heterobilayer $\rm{MoTe_{2}/WSe_{2}}$ system. The $U(1)$-slave spin method is used to study how the electronic band structure evolves as the interaction effect gradually increases [64]. We find a topological Kondo semimetal phase.
@@ II Model and method
where $\langle$.$\rangle$ . ⟩ denotes the expectation value, which is taken to be translationally invariant. We refer to the Appendix of Ref. [9] about the convention and further details of the method as well as on the procedure for numerical solution.
@@ III.1 Kondo semimetal
Figure 3: (a) The single valley dispersion relation of the renormalized bands along the high symmetry lines of the Brillouin zone in the Kondo semimetal phase. The color indicates the ratio of the orbital components. (b) Fermi surface contours for $\mathcal{K}$ valley in this Kondo semimetal phase. The blue and red lines stand for electron and hole pockets from the upper and lower bands, respectively. Both (a) and (b) are for the $1+1$ commensurate filling, with the parameters that result in an orbital-selective quasiparticle weight $Z_{\rm{Mo}}=0.16$ and $Z_{\rm{W}}=0.76$; $\epsilon_{D}=0$.
@@
The evolution of the quasiparticle weights for both orbitals versus the potential difference $\epsilon_{D}$ is shown in Fig. 2(b). Clearly, a large quasiparticle renormalization is observed in the $\rm{Mo}$ orbital, in particular when $\epsilon_{D}$ is close to zero. The drop in the quasiparticle weight in the heavy orbital is also observed in the correlated metal region when the total filling $1+x$ is close to $1$ (i.e., when $x$ is small) [9]. However, unlike that regime, where the conduction electron is barely renormalized ($Z_{W}\approx 1$), in the $\nu=1+1$ case, the quasiparticle weight number $Z_{W}\approx 0.8$. This is because the electron filling of the wide $c$ band is close to $1$, making its correlation effect more pronounced.
Fig. 2(c) shows the coherence temperature. It is estimated by taking both the hybridization and the residual intra-orbital hopping into account [9]. Specifically, $T_{\rm coh}\sim Z_{\rm Mo}D_{\rm Mo}+\frac{Z_{\rm Mo}Z_{\rm W}t^{2}_{cd}}{D_{\rm W}}$, which can be much higher than the usual estimation of the coherence temperature from the hybridization effect alone [65, 66, 67].
@@
The renormalized band structure along the high symmetry cut is shown in Fig. 3(a), where one can see that the electron and hole pockets originate from the upper and lower bands, respectively. As a consequence, the Fermi surface, which is shown in Fig. 3(b), comprises both electron and hole pockets.
@@ III.3 Spontaneous Hall Effect
Thus, we propose measurement of the spontaneous Hall effect as an experimental means of probing this topological semimetal phase. Our results make it clear that probing both the DC and AC responses will be instructive.
@@ IV Discussion
Several remarks are in order. Firstly, we have shown that the Kondo effect and band topology cooperate to generate a topological Kondo semimetal phase at the commensurate filling $\nu_{\rm Mo}=\nu_{\rm W}=1$. The topology is captured by the valley Berry curvature, which, in our case, is associated with the development of the flat band from the Kondo effect. This is analogous to what happens in the Weyl Kondo semimetal [35, 36, 37, 38, 39].
Experimentally, for the AB-stacked heterobilayer $\rm{MoTe_{2}/WSe_{2}}$, we note that the case when the total hole concentration $\nu_{\rm tot}=2$ has also been studied, providing evidence for the quantum spin Hall effect [4]. However, in the studied parameter region, the value of the displacement field is more likely to correspond to the situation with $\nu_{\rm Mo}+\nu_{\rm W}=2+0$. In this scenario, the electrons mostly reside in one of the layers, which is far away from the Mott insulating region, resulting in a band insulator. Our work applies to the case of a larger displacement field in the region of $\nu_{\rm Mo}+\nu_{\rm W}=1+1$.
Secondly, the analogue between the topological Kondo semimetals we discuss here and the notion of the Weyl Kondo semimetal [35, 37, 38, 39] may be even more direct. If the TMDC bilayers are stacked into a three-dimensional structure, the Kondo driven flat bands can develop Weyl nodes. The limitation placed by the separation of the two moiré bands may be overcome using the notion of emergent flat bands, in which the correlation-induced flat band is pinned to the Fermi energy even though the bare flat band is located away from it [62].
@@
Finally, the more localized $d$ orbitals will experience RKKY interactions. These interactions are expected to cause competition against the tendency towards forming the Kondo-driven flat bands. The expected global phase diagram in the commensurate ($1+1$) electron filling regime has been considered in the canonical Kondo lattice model systems [33, 79, 80]. As illustrated in Fig. 1(a), the competing phases involve certain symmetry-breaking order and its symmetry-preserving analogue. Theoretically, a systematic analysis of this competition could be realized through the extended dynamical mean field theory [81]; such a study is left for the future. Experimentally, it would be instructive to explore the Kondo-destruction phenomena  [43, 44, 59] in this commensurate filling setting.
@@ V Summary
Note added: After the completion of this manuscript and just before its posting on the arXiv, two independent works appeared on the arXiv [82, 83], in which related physics is studied in similar models for AB-stacked $\rm{MoTe_{2}/WSe_{2}}$ though the spontaneous Hall effect is not considered.
@@ Acknowledgements.
We thank Jennifer Cano, Kin Fai Mak, Silke Paschen, Jie Shan and Wenjin Zhao for helpful discussions. This work has primarily been supported by the NSF Grant No. DMR-2220603 (model construction, F.X. and L.C.), by the Air Force Office of Scientific Research under Grant No. FA9550-21-1-0356 (conceptualization and model calculation, F.X., L.C. and Y.F.), and by the Robert A. Welch Foundation Grant No. C-1411 and the Vannevar Bush Faculty Fellowship No. ONR-VB N00014-23-1-2870 (Q.S.). The majority of the computational calculations have been performed on the Shared University Grid at Rice funded by NSF under Grant No. EIA-0216467, a partnership between Rice University, Sun Microsystems, and Sigma Solutions, Inc., the Big-Data Private-Cloud Research Cyberinfrastructure MRI-award funded by NSF under Grant No. CNS-1338099, and the Extreme Science and Engineering Discovery Environment (XSEDE) by NSF under Grant No. DMR170109. Q.S. acknowledges the hospitality of the Aspen Center for Physics, which is supported by NSF grant No. PHY-2210452.

### v2

@@ I Introduction
Figure 1: (a) Global phase diagram of Kondo lattice models at commensurate filling factor with ($\nu_{c}\approx\nu_{d}\approx 1$), where $c$ and $d$ correspond to light and heavy orbitals, respectively. Vertical axis $G$ stands for the frustration strength, and horizontal axis $\rho J_{K}$ describes the Kondo coupling. The blue and black lines correspond to Kondo-destruction (dehybridization) and magnetic transitions, respectively. They separate the parameter region into four phases: (i) $M_{S}$, a magnetic ordered phase with a small Fermi surface; (ii) $P_{S}$, a paramagnetic phase with a small Fermi surface; (iii) $M_{L}$, a magnetic ordered phase with a large Fermi surface; and (iv) Kondo-screened phase, which is either a Kondo semimetal (KSM) or a Kondo insulator (KI). The dashed lines describe distinct sequences of quantum phase transitions between the Kondo-destroyed phases and the Kondo-screened ones. Adapted from Ref. [33]. (b) The lattice structure of the $2H$ phase of single layer transition metal dichalcogenides. The purple and yellow orbs represent the metal and chalcogen atoms, respectively. This lattice structure already breaks the inversion symmetry. Lattice structure data retrieved from the Materials Project for Te2Mo (mp-1023935) from database version v2023.11.1 [34] (c) Spin-valley locking in the TMDC materials. $\mathcal{K},\mathcal{K}^{\prime}$ denote the two valleys in the TMDC system. Due to the AB-stacking structure, the spin polarizations from the two TMDC layers in the same valley are different, thereby reducing the interlayer hybridization. (d) Noninteracting band structure of the two-orbital effective tight binding model with $t_{c}=9\leavevmode\nobreak\ \rm meV$, $t_{d}=4.5\leavevmode\nobreak\ \rm meV$, $t_{cd}=1.5\leavevmode\nobreak\ \rm meV$ and $\epsilon_{D}=0\leavevmode\nobreak\ \rm meV$, which well-captures the properties of the active moiré bands in AB-stacked $\rm MoTe_{2}/WSe_{2}$. The definition of the high-symmetry points and high-symmetry lines are also shown.
@@ I.2 Kondo effect and its realization in heterobilayer TMDCs
Kondo lattice systems [42] represent a canonical setting to realize a landscape of correlated quantum phases. The contrast in the width of the $f$-electron band and that of the $spd$-electron bands causes orbital-selective Mott correlations [43, 44]. In particular, metallic quantum criticality and strange metallicity have been the main stay of the modern era of this field, fueled by the notion of Kondo destruction [45, 46, 47, 48]. This notion captures the dynamical competition between the Kondo and Ruderman–Kittel–Kasuya–Yosida (RKKY) interactions, leading, for example, to an interacting quantum critical point with dynamical “Planckian” scaling [46]. The Kondo destruction notion is supported by a large body of experiments in a variety of quantum critical heavy fermion metals  [49, 50, 51, 52, 53, 54, 55]. When combined with geometrical frustration and related quantum fluctuations of the local moments, this notion leads to a global phase diagram [56, 57, 58, 59], delineating the interplay between the local-moment magnetism and paramagnetism, the “large” Fermi surface associated with the Kondo-screened ground state, and the “small” Fermi surface associated with the destruction of the Kondo effect [60, 61].
@@
In other words, if we compare the energy scale in the AB-stacked heterobilayer TMDC system, it will fall into the parameter regime with $t_{c}>t_{d}>\tilde{t}_{cd}$ (where $\tilde{t}_{cd}\approx\pi\rho_{0,c}t^{2}_{cd}$ with $\rho_{0,c}$ being the typical bare density of states of the $c$ band). This is different from what happens in the canonical heavy fermion systems [42, 43, 44] or in the effective Kondo lattice models realized in geometrically-enabled-flat-band systems [62, 63, 64], where $t_{c}>\tilde{t}_{cd}>t_{d}$. Nonetheless, the Coulomb interaction can significantly amplify the renormalized bandwidths difference between the $d$ and $c$ orbitals, with the $d$ orbital approaching the local spin moment limit. Additionally, the presence of hybridization $\tilde{t}_{cd}$, combined with the fact that the $d$ orbitals form a triangular lattice with geometric frustration, makes this system a potential simulator of Kondo physics.
We also note that the “inverted” hierarchy of energy scales has important consequences. Such a large value of $t_{d}$ obviously leads to strong local charge fluctuations, which means that the Kondo-driven correlated phases are in the so-called “mixed valence” regime. For brevity, it is still usually referred to as Kondo insulators or Kondo semimetals in the literature, as the Kondo effect and Kondo destruction phase transition can still be well-defined in this regime [33, 65]. Another concerns the coherent temperature scale, which characterizes the crossover between the high-temperature incoherent scattering regime and the low temperature Fermi liquid regime. Due to the inverted energy scale hierarchy, the coherence temperature relevant to the strongly correlated metal region depends not only on the hybridization but also on the contribution from the intra-orbital kinetic energy of the heavy electron itself [9]. Accordingly, this coherent temperature scale will be unexpectedly large, as indeed observed experimentally [3]. As we will show, this energy scale hierarchy plays an important role in realizing a topological Kondo semimetal phase in the commensurate filling that we will study.
@@ I.3 Topological Kondo lattice Systems at commensurate fillings: Topological Kondo semimetals
In recent years, topological Kondo lattice models [35, 36, 37] and materials [38, 39] have been advanced, which realize the cooperation between the inherent topology of the noninteracting bands and the correlation physics associated with the Kondo effect at commensurate electron fillings. At such fillings, a Kondo insulator gap would have been opened [66, 67]; its prevention by the the lattice symmetry that underlies the band topology gives rise to Weyl Kondo semimetals.
@@ I.4 Scope of the present work
Recently, the Kondo effect and its destruction have been theoretically discussed in models for the AB-stacked $\rm{MoTe_{2}/WSe_{2}}$ [8, 9]. The inherently topological nature of the electronic bands raises the prospect that this moiré structure can emulate topological Kondo lattice systems. It is the objective of this work to explore this prospect.
In our previous work [9], we focused on incommensurate fillings $\nu_{\rm tot}=1+x$ of this $\rm{MoTe_{2}/WSe_{2}}$ hetero-bilayer system, especially in the regime with rather small $x$. As we have emphasized in Sec. I.3, the commensurate filling case $\nu_{\rm tot}=1+1$ is always considered to be special in the context of heavy fermion systems, which corresponds to potential Kondo topological semimetal or Kondo insulator phases, instead of metallic phases such as heavy Fermi liquid. In this work, the $U(1)$-slave spin method is used to study how the electronic band structure evolves as the interaction effect gradually increases [68]. We find a topological Kondo semimetal phase.
@@ II Model and method
[…]
@@ III.1 Kondo semimetal
Figure 3: (a) The single valley dispersion relation of the renormalized bands along the high symmetry lines of the Brillouin zone in the Kondo semimetal phase. The color indicates the ratio of the orbital components. (b) Fermi surface contours for $\mathcal{K}$ valley in this Kondo semimetal phase. The blue and red lines stand for electron and hole pockets from the upper and lower bands, respectively. Both (a) and (b) are for the $1+1$ commensurate filling, with the parameters that result in an orbital-selective quasiparticle weight $Z_{\rm{Mo}}=0.16$ and $Z_{\rm{W}}=0.76$; $\epsilon_{D}=0\,\rm meV$.
@@
The evolution of the quasiparticle weights for both orbitals versus the potential difference $\epsilon_{D}$ is shown in Fig. 2(b). Clearly, a large quasiparticle renormalization is observed in the $\rm{Mo}$ orbital, in particular in the parameter regime highlighted by the light blue shaded region, where $\epsilon_{D}$ is close to zero. The drop in the quasiparticle weight in the heavy orbital is also observed in the correlated metal region when the total filling $1+x$ is close to $1$ (i.e., when $x$ is small) [9]. However, unlike that regime, where the conduction electron is barely renormalized ($Z_{W}\approx 1$), in the $\nu=1+1$ case, the quasiparticle weight number $Z_{W}\approx 0.8$. This is because the electron filling of the wide $c$ band is close to $1$, making its correlation effect more pronounced.
Fig. 2(c) shows the coherence temperature. It is estimated by taking both the hybridization and the residual intra-orbital hopping into account [9]. Specifically, the coherence temperature can be estimated through $T_{\rm coh}\sim Z_{\rm Mo}D_{\rm Mo}+\frac{Z_{\rm Mo}Z_{\rm W}t^{2}_{cd}}{D_{\rm W}}$, in which $D_{\rm Mo}\sim 20\rm\,meV$ and $D_{\rm W}\sim 40\rm\,meV$ are the half bandwidths of the two active bands. This value can be much higher than the usual estimation of the coherence temperature from the hybridization effect alone [69, 70, 71].
@@
The renormalized band structure at $\epsilon_{D}=0\rm\,meV$ along the high symmetry cut is shown in Fig. 3(a), where one can see that the electron and hole pockets originate from the upper and lower bands, respectively. As a consequence, the Fermi surface, which is shown in Fig. 3(b), comprises both electron and hole pockets. We also argue that the qualitative features of the electronic structure is not very sensitive to the value of displacement field strength within this light blue shaded region. Therefore, in the following, we will mostly use the electronic structure at $\epsilon_{D}=0\rm\,meV$ as a representative case. In Appendix B, we present several renormalized band structure plots at different displacement field potential strength.
@@ III.3 Spontaneous Hall Effect
Thus, we propose measurement of the spontaneous Hall effect as an experimental means of probing this topological semimetal phase. Our results make it clear that probing both the DC and AC responses will be instructive. The Berry curvature hexaople can be as large as $Q_{xxxxy}\sim 10^{3}a_{M}^{3}$, where $a_{M}$ is the moire lattice scale. The large value of $a_{M}$ enhances the prospect of experimentally detecting this nonlinear response. Because it is comparable to the length scale of the nonlinear response $\ell_{E}=\hbar/(eE\tau)\sim 10^{-7}\,{\rm m}$ for electric field strength $E_{x}\sim 10\,{\rm V\cdot mm^{-1}}$ and relaxation time $\tau\sim 10^{-12}\,{\rm s}$. However, the exact value of $Q_{xxxxy}$ is sensitive to whether the Fermi surface overlap with the peak of Berry curvature, and whether it can be measured under current experimental capability depends on the detailed band structure, which is beyond the scope of our model study.
@@ IV Discussion
Several remarks are in order. First, we have shown that the Kondo effect and band topology cooperate to generate a topological Kondo semimetal phase at the commensurate filling $\nu_{\rm Mo}=\nu_{\rm W}=1$. The topology is captured by the valley Berry curvature, which, in our case, is associated with the development of the flat band from the Kondo effect. This is analogous to what happens in the Weyl Kondo semimetal [35, 36, 37, 38, 39].
Experimentally, for the AB-stacked heterobilayer $\rm{MoTe_{2}/WSe_{2}}$, we note that the case when the total hole concentration $\nu_{\rm tot}=2$ has also been studied, providing evidence for the quantum spin Hall effect [4]. However, in the studied parameter region, the value of the displacement field is more likely to correspond to the situation with $\nu_{\rm Mo}+\nu_{\rm W}=2+0$. In this scenario, the electrons mostly reside in one of the layers, which is far away from the Mott insulating region, resulting in a band insulator. Our work applies to the case of a larger displacement field in the region of $\nu_{\rm Mo}+\nu_{\rm W}=1+1$. We note that two independent recent works [78, 79] have explored related physics in similar models for AB-stacked MoTe2/WSe2, though the spontaneous Hall effect is not considered there.
Second, the analogue between the topological Kondo semimetals we discuss here and the notion of the Weyl Kondo semimetal [35, 37, 38, 39] may be even more direct. If the TMDC bilayers are stacked into a three-dimensional structure, the Kondo driven flat bands can develop Weyl nodes. The limitation placed by the separation of the two moiré bands may be overcome using the notion of emergent flat bands, in which the correlation-induced flat band is pinned to the Fermi energy even though the bare flat band is located away from it [63].
@@
Third, the more localized $d$ orbitals will experience RKKY interactions. These interactions are expected to cause competition against the tendency towards forming the Kondo-driven flat bands. The expected global phase diagram in the commensurate ($1+1$) electron filling regime has been considered in the canonical Kondo lattice model systems [33, 85, 86]. As illustrated in Fig. 1(a), the competing phases involve certain symmetry-breaking order and its symmetry-preserving analogue. Theoretically, a systematic analysis of this competition could be realized through the extended dynamical mean field theory [87]; such a study is left for the future. Experimentally, it would be instructive to explore the Kondo-destruction phenomena  [43, 44, 60] in this commensurate filling setting.
Finally, our discussion are based on the assumption that there are no magnetic orders or density wave orders with spontaneous translation or time-reversal symmetry breaking. Although there is currently no experimental evidence for the existence of such states in the relevant parameter regimes, their presence could indeed affect the potential topological responses. A detailed study of these possibilities lies beyond the scope of the present study based on parton construction, and would be an interesting direction for future work.
@@ Acknowledgements.
We thank Jennifer Cano, Kin Fai Mak, Silke Paschen, Jie Shan and Wenjin Zhao for helpful discussions. This work has primarily been supported by the NSF Grant No. DMR-2220603 (model construction, F.X. and L.C.), by the Air Force Office of Scientific Research under Grant No. FA9550-21-1-0356 (conceptualization and model calculation, F.X., L.C. and Y.F.), and by the Robert A. Welch Foundation Grant No. C-1411 and the Vannevar Bush Faculty Fellowship No. ONR-VB N00014-23-1-2870 (Q.S.). The majority of the computational calculations have been performed on the Shared University Grid at Rice funded by NSF under Grant No. EIA-0216467, a partnership between Rice University, Sun Microsystems, and Sigma Solutions, Inc., the Big-Data Private-Cloud Research Cyberinfrastructure MRI-award funded by NSF under Grant No. CNS-1338099, and the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) by NSF under Grant No. DMR170109. Q.S. acknowledges the hospitality of the Aspen Center for Physics, which is supported by NSF grant No. PHY-2210452. The data that support the findings of this article are openly available [88].
Appendix A $U(1)$ slave spin approach
In this appendix, we briefly review the method of $U(1)$ slave-spin approach. A physical fermion operator $\alpha_{i\alpha\tau}^{\dagger}$ ($\alpha$ can be $c$ or $d$) is mapped to a fermionic operator $f^{\dagger}_{i\alpha\tau}$ and a spin operator $o^{\dagger}_{i\alpha\tau}$:
$\alpha^{\dagger}_{i\alpha\tau}\rightarrow f^{\dagger}_{i\alpha\tau}o^{\dagger}_{i\alpha\tau}\,,$ (5)
in which $i$, $\alpha$ and $\tau$ stand for unit cell, orbital and valley indices. The slave spin operator has the following form:
$o^{\dagger}_{i\alpha\tau}=P^{+}_{i\alpha\tau}S^{+}_{i\alpha\tau}P^{-}_{i\alpha\tau}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ P^{\pm}_{i\alpha\tau}=\frac{1}{\sqrt{\frac{1}{2}\pm S^{z}_{i\alpha\tau}}}\,,$ (6)
where $S^{\pm}$ and $S^{z}$ operate on the space of spin-$\frac{1}{2}$. We note that such parton construction enlarges the local Hilbert space. Among the four states spanned by the slave spin and slave fermion, only two of them correspond to the physical states:
$|n_{i\alpha\tau}=1\rangle\leftrightarrow$ $|n^{f}_{i\alpha\tau}=1\rangle|S^{z}_{i\alpha\tau}=\uparrow\rangle\,,$ (7)
$|n_{i\alpha\tau}=0\rangle\leftrightarrow$ $|n^{f}_{i\alpha\tau}=0\rangle|S^{z}_{i\alpha\tau}=\downarrow\rangle\,,$ (8)
[…]
Obviously, the local constraint is given by $\langle S^{z}_{i\alpha\tau}\rangle+1/2=\langle n^{f}_{i\alpha\tau}\rangle$. On the saddle-point level, such constraint can be employed by adding an extra Lagrange multiplier term into the parton Hamiltonian:
$H_{\lambda}=\sum_{i\alpha\tau}\lambda_{i\alpha\tau}\left(S^{z}_{i\alpha\tau}+\frac{1}{2}-n^{f}_{i\alpha\tau}\right)\,,$ (11)
Using the parton operators to rewrite the Hamiltonian and and applying a mean-field decoupling between the slave fermions and slave spins, the Hamiltonian can be separated into two distinct parts $H^{f}$ and $H^{S}$. In particular, the slave-spin part becomes local terms with the following form:
$H^{S}=H_{I}^{S}+\sum_{\alpha\tau}\lambda_{\alpha}S^{z}_{\alpha\tau}+\sum_{\alpha\tau}\left[h_{\alpha}\frac{S^{+}_{\alpha\tau}}{\sqrt{n^{f}_{\alpha}(1-n^{f}_{\alpha})}}+{\rm h.c.}\right]\,,$ (12)
in which $H_{I}^{S}$ is defined in Eq. (II) in the main text, which stands for the “on-site” part of the interaction written using the slave-spin operators and $n^{f}_{\alpha}=\langle f^{\dagger}_{i\alpha\tau}f_{i\alpha\tau}\rangle$ is the fermion density expectation value of the orbital $\alpha$. Here we have assumed that the parameters $\lambda$, $h_{\alpha}$ and $Z$ are independent on unit cell index $i$ and valley index $\tau$, which represents a non-magnetic state without translation symmetry breaking orders. The quasiparticle weights are defined as the expectation value of the slave spin operators in the following way:
$Z_{\alpha}=\frac{\langle S^{+}_{\alpha\tau}\rangle}{\sqrt{n^{f}_{\alpha}(1-n^{f}_{\alpha})}}\,,$ (13)
and the “bath” fields are defined as:
$h_{c}=$ $\frac{1}{N}\sum_{\mathbf{k}}\epsilon_{cd}(\mathbf{k})\sqrt{Z_{d}}\langle f^{\dagger}_{\mathbf{k}c\tau}f_{\mathbf{k}d\tau}\rangle$
$+\frac{1}{N}\sum_{\mathbf{k}}\epsilon_{c}(\mathbf{k})\sqrt{Z_{c}}\langle f^{\dagger}_{\mathbf{k}c\tau}f_{\mathbf{k}c\tau}\rangle\,,$ (14)
$h_{d}=$ $\frac{1}{N}\sum_{\mathbf{k}}\epsilon_{cd}(\mathbf{k})\sqrt{Z_{c}}\langle f^{\dagger}_{\mathbf{k}d\tau}f_{\mathbf{k}c\tau}\rangle$
$+\frac{1}{N}\sum_{\mathbf{k}}\epsilon_{d}(\mathbf{k})\sqrt{Z_{d}}\langle f^{\dagger}_{\mathbf{k}d\tau}f_{\mathbf{k}d\tau}\rangle\,,$ (15)
in which $\epsilon_{c}(\mathbf{k})$, $\epsilon_{d}(\mathbf{k})$ and $\epsilon_{cd}(\mathbf{k})$ stands for the Fourier transformations of the intralayer and interlayer hopping terms, respectively. On the fermionic side, the mean-field Hamiltonian has the following quadratic form:
$H^{f}=$ $\sum_{\langle\langle ij\rangle\rangle,\tau}t_{c}Z_{c}e^{\tau i\phi^{c}_{ij}}f^{\dagger}_{ic\tau}f_{jc\tau}+\sum_{\langle\langle ij\rangle\rangle,\tau}t_{d}Z_{d}e^{\tau i\phi^{d}_{ij}}f^{\dagger}_{id\tau}f_{jd\tau}$
$+\sum_{\langle ij\rangle,\tau}\left(t_{cd}\sqrt{Z_{c}Z_{d}}f^{\dagger}_{ic\tau}f_{jd\tau}+{\rm h.c.}\right)$
$+\sum_{i\tau}f^{\dagger}_{ic\tau}f_{ic\tau}\left(\frac{\epsilon_{D}}{2}-\lambda_{c}+\lambda_{c}^{0}\right)$
$+\sum_{i\tau}f^{\dagger}_{id\tau}f_{id\tau}\left(-\frac{\epsilon_{D}}{2}-\lambda_{d}+\lambda_{d}^{0}\right)\,,$ (16)
and the parameters $\lambda^{0}_{\alpha}$ are defined as:
$\lambda_{\alpha}^{0}=-\sqrt{Z_{\alpha}}|h_{\alpha}|\frac{2n_{\alpha}^{f}-1}{n^{f}_{\alpha}(1-n^{f}_{\alpha})}\,.$ (17)
These terms guarantee the quasiparticle weights of both orbitals will be reduced to $1$ when the interaction terms are turned off.
We note that the quasiparticle weights $Z_{\alpha}$, which appear in the fermion Hamiltonian $H^{f}$ as parameters, can be determined from the ground state of the spin Hamiltonian $H^{S}$. Similarly, the bath fields $h_{\alpha}$, which appear in the spin Hamiltonian $H^{S}$ as parameters, can be determined from the ground state of the fermion Hamiltonian $H^{f}$. Additionally, the Lagrange multipliers $\lambda_{\alpha}$ are solved from the local constraint equation. Hence, with all these equations taken into account, the parameters $\lambda_{\alpha}$, $\lambda_{\alpha}^{0}$ and $Z_{\alpha}$ can be determined self-consistently.
Appendix B Electronic structure under different displacement field strengths
Figure 6: (a-b) The renormalized band structure and the Fermi surface shape in a single valley with displacement field potential strength $\epsilon_{D}=0\,\rm meV$. (c-d) The renormalized band structure and the Fermi surface in a single valley with displacement field potential strength $\epsilon_{D}=-6\rm\,meV$. (e-f) The single value renormalized band structure and the Fermi surface under displacement field potential strength $\epsilon_{D}=6\rm\,meV$. Subfigures (a-b) are identical to Fig. 3 for easy comparison.
In this appendix, we provide the numerical results of the electronic structure, including the band structure and the Fermi surface structure at the commensurate filling under different displacement field strength. As shown in Fig. 2, the quasiparticle weight of the Mo orbital is strongly suppressed in the light blue shaded region. In Figs. 3, 4 and 5, we show the Berry curvature and its dipole moment at a displacement field strength of $\epsilon_{D}=0,\rm meV$, which lies within this interval.
Here we present the renormalized band structure and Fermi surface shape at different displacement field potential strength. In Figs. 6(c-f), the electronic structure at displacement field potential strength $\epsilon_{D}=-6\,\rm meV$ and $\epsilon_{D}=6\,\rm meV$ are shown. By comparing these results with the band structure and Fermi surface at $\epsilon_{D}=0\,\rm meV$ [shown in Fig. 6(a-b)], we find that the qualitative features, such as the amount of Fermi pockets, remain largely unaffected by the displacement field strength. However, some quantitative differences can be observed. For example, due to the slightly different quasiparticle weight of the Mo orbital, its corresponding bandwidth is also slightly narrower at $\epsilon_{D}=6\,\rm meV$ than at $\epsilon_{D}=0\,\rm meV$.